if the mass of earth is 100 times the mass of moon & its diameter 5 times that of moon compare the weight on surface of moon with that of earth
 

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Weight of body
W = G*m * m(b)/r²
Now
We/Wm = [100*m(m)* m(b)/25 R²m] / [m(m)* m(b)/R²(m)]
We/Wm = 100/25 = 4
 
Regards
Arun (askIITians forum expert)